Write an equation of the line that passes through the point and has the given slope. Write the equation in slope-intercept form (y= mx + b)
(3,4), m = 1/2


Answer by SoOoDumb(14)

x,y
(3,4)
m=1/2
y=mx+b
4=1/2(3)+b
4=1.5+b
2.5=b
y=1/2x+2.5